OpenStudy (anonymous):

For the given functions f and g, find the requested composite function value. f(x) = 2x + 4, g(x) = 4x2 + 3; Find (g * f)(3).

OpenStudy (mathmate):

gof(3)=g(f(3))=g(2*3+4)=g(10)=4(10)^2+4=404